# Search for spatial coincidences between galaxy mergers and Fermi-LAT 4FGL-DR4 sources

### v1

@@ III Results
Next, we applied the matching criterion defined in Equation 2 to our primary sample of 3,166 morphologically selected galaxy mergers identified from SDSS imaging. This resulted in identification of 38 spatial coincidences between Fermi 4FGL-DR4 sources and galaxy mergers from our sample. To assess the likelihood of true physical associations versus chance alignments, we computed Poisson-based match probabilities for each pair using Eq. 2. Among these 38 matches, 21 exhibit p-values below 0.05, indicating a low probability of random association. We classify these as statistically secure matches. According to source classifications in the 4FGL-DR4 catalog, the 21 matches within the 4 ellipses include four Flat Spectrum Radio Quasars (FSRQs), six BL Lacertae objects (BLL), three Blazar Candidates of Uncertain type (BCUs), and three Radio Galaxies (RDGs). The remaining five are unassociated sources (UIDs) lacking confirmed multiwavelength counterparts. Notably, none of the secure matches is classified as a pulsar. The five UIDs span gamma-ray flux densities in the range , suggesting the possible existence of a previously unrecognized population of gamma-ray sources associated with galaxy mergers.
@@
Overall, our analysis reveals that approximately 0.67% (21 out of 3,166) of galaxy mergers exhibit statistically robust associations with gamma-ray sources. The predominance of blazar-type objects among these secure matches accounting for 62% is consistent with their known dominance in the extragalactic gamma-ray sky. However, the presence of radio galaxies and a significant fraction of unassociated sources (24%) suggests that merger-driven gamma-ray activity may originate from a broader range of physical mechanisms than previously recognized. To visualize these associations, Figure 2 presents a multi-panel view of Fermi sources with at least one galaxy merger candidate within the corresponding positional uncertainty ellipse. In each panel, the ellipse denotes the Fermi positional uncertainty region, the blue dot marks the Fermi source position, and the red cross indicates the galaxy merger counterpart with the lowest p-value. In Table 1 we depict all the secure matches with their coordinates, SDSS -mag, spectroscopic redshifts, flux densities, p-values, angular separation and classification. To explore marginal cases, we extended the analysis to include matches with , identifying 24 potential associations. Of these, 21 are already included in the secure sample (). The remaining three additional sources are 4FGL J1454.0+4927 (BCU), 4FGL J1523.2+2818 (BCU), and 4FGL J1223.3+1213 (BLL), with gamma-ray flux densities of , , and , and corresponding p-values of 0.069, 0.071, and 0.076, respectively. These sources are spatially associated with the SDSS galaxies SDSS J145506.73+493614.7, SDSS J152228.74+281402.9, and SDSS J122256.10+121739.2, which have spectroscopic redshifts of 0.0945, 0.0715, and 0.0905, respectively.
IV Conclusions
@@
We have presented the first dedicated cross-matching analysis between two galaxy merger catalogs and the Fermi-LAT 4FGL-DR4 gamma-ray source catalog Ballet et al. (2023). The first catalog consists of a sample of 70 galaxy pairs from the Canadian Network for Observational Cosmology Field Galaxy Redshift Survey and found no statistically significant associations within the 4 positional uncertainty threshold (). The second catalog comprises of a high-confidence sample of 3,166 morphologically selected galaxy mergers Ackermann et al. (2018) and the Fermi-LAT 4FGL-DR4 gamma-ray source catalog Ballet et al. (2023). Using a conservative 4 positional uncertainty criterion and a Poisson-based statistical framework, we identified 38 spatial coincidences, of which 21 exhibit p-values below 0.05. These matches predominantly include well-known gamma-ray source classes, FSRQs, BL Lacs, and radio galaxies, supporting the hypothesis that merger-driven processes can fuel AGN activity and high-energy emission. Notably, five secure associations are with unclassified (unassociated) Fermi sources from the fourth FGL-DR4 catalog, suggesting that merging systems may host previously unrecognized gamma-ray populations.
@@
While blazar-like AGN dominate the matched sample, the presence of radio galaxies, UIDs and the broad range of observed gamma-ray flux densities point to a diversity of emission mechanisms. These may include merger-induced starbursts, large-scale shocks, or obscured AGN activity, highlighting the need for deeper multi-wavelength follow-up Torres (2004); Ackermann et al. (2012). The consistent appearance of unassociated sources within the secure sample underscores the value of galaxy merger catalogs as informative priors for gamma-ray source identification. As gamma-ray localization continues to improve and future optical surveys such as Vera Rubin LSST, Euclid, and DESI expand the census of galaxy mergers, the ability to resolve such associations will significantly increase. We shall also conduct a targeted search for some specific merger candidates, such as UGC 12914/5, UGC 813/6, and VV 114 in a forthcoming work.
@@
4FGL J0016.20016 | 3.9988 | 0.3049 | 16.498 | 0.0396 | 0.0081 | 2.50 | FSRQ | |
4FGL J0813.4+2531 | 122.8163 | 25.1794 | 14.146 | 0.0137 | 0.0113 | 3.99 | FSRQ | |
4FGL J0815.6+3641 | 123.8059 | 36.7766 | 16.851 | 0.0420 | 0.0252 | 3.77 | FSRQ | |
4FGL J0844.5+3035 | 131.0952 | 30.5990 | 15.716 | 0.0748 | 0.0042 | 1.65 | UID | |
4FGL J1008.0+0028 | 151.9620 | 0.5303 | 17.075 | 0.0934 | 0.0111 | 3.13 | BLL | |
4FGL J1018.9+1043 | 154.6662 | 10.7793 | 14.790 | 0.0513 | 0.0058 | 0.53 | BCU | |
4FGL J1105.8+3944 | 166.4743 | 39.7825 | 15.958 | 0.0991 | 0.0026 | 2.26 | UID | |
4FGL J1112.0+2607 | 168.0185 | 26.0447 | 17.046 | 0.0696 | 0.0182 | 2.64 | UID | |
4FGL J1116.6+2915 | 169.1460 | 29.2497 | 14.657 | 0.0500 | 0.0006 | 0.80 | RDG | |
4FGL J1144.9+1937 | 176.2744 | 19.6365 | 17.479 | 0.0162 | 0.0083 | 1.97 | RDG | |
4FGL J1145.50340 | 176.3724 | 3.6637 | 16.811 | 0.0790 | 0.0122 | 0.83 | BLL | |
4FGL J1202.9+5141 | 180.7901 | 51.6782 | 15.849 | 0.0607 | 0.0060 | 1.67 | BCU | |
4FGL J1245.8+0232 | 191.5785 | 2.6659 | 17.394 | 0.0811 | 0.0344 | 3.31 | BLL | |
4FGL J1353.2+3740 | 208.3187 | 37.7076 | 16.675 | 0.0630 | 0.0076 | 1.26 | BLL | |
4FGL J1411.8+5249 | 212.9420 | 52.7960 | 16.248 | 0.0766 | 0.0028 | 2.64 | BLL | |
4FGL J1435.4+3338 | 218.7278 | 33.8222 | 15.394 | 0.0586 | 0.0171 | 3.00 | UID | |
4FGL J1441.7+1836 | 220.5586 | 18.6907 | 15.233 | 0.0710 | 0.0035 | 3.91 | BCU | |
4FGL J1508.8+2708 | 227.1547 | 27.1589 | 15.109 | 0.0760 | 0.0005 | 3.41 | BLL | |
4FGL J1518.6+0614 | 229.5286 | 6.3552 | 16.892 | 0.0795 | 0.0274 | 2.68 | RDG | |
4FGL J1600.4+0407 | 240.2073 | 4.1848 | 17.213 | 0.0877 | 0.0137 | 3.42 | UID | |
4FGL J1606.5+2717 | 241.6173 | 27.1373 | 17.119 | 0.0458 | 0.0480 | 3.70 | FSRQ |
Acknowledgements.
@@
SM gratefully acknowledges the Ministry of Education (MoE), Government of India, for their consistent support through the research fellowship, which has been instrumental in facilitating the successful completion of this work. The authors also thank Lizhong Zhang for useful correspondence.References
- Atwood et al. (2009) W. B. Atwood, A. A. Abdo, M. Ackermann, W. Althouse, B. Anderson, M. Axelsson, L. Baldini, J. Ballet, D. L. Band, G. Barbiellini, et al., Astrophys. J. 697, 1071 (2009), eprint 0902.1089.
- Ballet et al. (2023) J. Ballet, P. Bruel, T. H. Burnett, B. Lott, and The Fermi-LAT collaboration, arXiv e-prints arXiv:2307.12546 (2023), eprint 2307.12546.
- Abdollahi et al. (2022) S. Abdollahi, F. Acero, L. Baldini, J. Ballet, D. Bastieri, R. Bellazzini, B. Berenji, A. Berretta, E. Bissaldi, R. D. Blandford, et al., Astrophys. J. Suppl. Ser. 260, 53 (2022), eprint 2201.11184.
- Mayer and Becker (2024) M. G. F. Mayer and W. Becker, Astron. & Astrophys. 684, A208 (2024), eprint 2401.17295.
- Ulgiati et al. (2025) A. Ulgiati, S. Paiano, F. Pintore, T. D. Russell, B. Sbarufatti, C. Pinto, E. Ambrosi, A. D’Aì, G. Cusumano, and M. Del Santo, Astron. & Astrophys. 694, A176 (2025), eprint 2412.19314.
- Bruzewski et al. (2023) S. Bruzewski, F. K. Schinzel, and G. B. Taylor, Astrophys. J. 943, 51 (2023), eprint 2212.05115.
- Zhang et al. (2022) L. Zhang, J. D. Vieira, M. Ajello, M. A. Malkan, M. A. Archipley, J. Capota, A. Foster, and G. Madejski, Astrophys. J. 939, 117 (2022), eprint 2206.06228.
- Ramakrishnan and Desai (2024) V. Ramakrishnan and S. Desai, arXiv e-prints arXiv:2412.03900 (2024), eprint 2412.03900.
- Negro et al. (2023) M. Negro, M. Crnogorčević, E. Burns, E. Charles, L. Marcotulli, and R. Caputo, Astrophys. J. 951, 83 (2023), eprint 2304.10934.
- Himes et al. (2025) M. Himes, P. Jagannathan, D. A. Frail, F. Schinzel, N. Gupta, S. A. Balashev, F. Combes, P. P. Deka, H. R. Klöckner, E. Momjian, et al., Astrophys. J. 980, 81 (2025), eprint 2412.12470.
- Hopkins et al. (2006) P. F. Hopkins, L. Hernquist, T. J. Cox, T. Di Matteo, B. Robertson, and V. Springel, Astrophys. J. Suppl. Ser. 163, 1 (2006), eprint astro-ph/0506398.
- Kaviraj (2025) S. Kaviraj, arXiv e-prints arXiv:2506.09136 (2025), eprint 2506.09136.
- Bouri et al. (2025) S. Bouri, P. Parashari, M. Das, and R. Laha, Phys. Rev. D 111, 063059 (2025), eprint 2404.06539.
- Lisenfeld and Völk (2010) U. Lisenfeld and H. J. Völk, Astron. & Astrophys. 524, A27 (2010), eprint 1009.1659.
- Kashiyama and Mészáros (2014) K. Kashiyama and P. Mészáros, Astrophys. J. Lett. 790, L14 (2014), eprint 1405.3262.
- Yuan et al. (2018) C. Yuan, P. Mészáros, K. Murase, and D. Jeong, Astrophys. J. 857, 50 (2018), eprint 1712.09754.
- Vorontsov-Velyaminov et al. (2001) B. A. Vorontsov-Velyaminov, R. I. Noskova, and V. P. Arkhipova, Astronomical and Astrophysical Transactions 20, 717 (2001).
- Ahumada et al. (2020) R. Ahumada, C. Allende Prieto, A. Almeida, F. Anders, S. F. Anderson, B. H. Andrews, B. Anguiano, R. Arcodia, E. Armengaud, M. Aubert, et al., Astrophys. J. Suppl. Ser. 249, 3 (2020), eprint 1912.02905.
- Ackermann et al. (2018) S. Ackermann, K. Schawinski, C. Zhang, A. K. Weigel, and M. D. Turp, MNRAS 479, 415 (2018), eprint 1805.10289.
- Patton et al. (2005) D. R. Patton, J. K. Grant, L. Simard, C. J. Pritchet, R. G. Carlberg, and K. D. Borne, Astron. J. 130, 2043 (2005), eprint astro-ph/0507417.
- Hwang and Chang (2009) C.-Y. Hwang and M.-Y. Chang, Astrophys. J. Suppl. Ser. 181, 233 (2009).
- Gimeno et al. (2004) G. N. Gimeno, R. J. Díaz, and G. J. Carranza, Astron. J. 128, 62 (2004).
- Taylor (2005) M. B. Taylor, in Astronomical Data Analysis Software and Systems XIV, edited by P. Shopbell, M. Britton, and R. Ebert (2005), vol. 347 of Astronomical Society of the Pacific Conference Series, p. 29.
- Browne and Cohen (1978) I. W. A. Browne and A. M. Cohen, MNRAS 182, 181 (1978).
- Downes et al. (1986) A. J. B. Downes, J. A. Peacock, A. Savage, and D. R. Carrie, MNRAS 218, 31 (1986).
- Biggs et al. (2011) A. D. Biggs, R. J. Ivison, E. Ibar, J. L. Wardlow, H. Dannerbauer, I. Smail, F. Walter, A. Weiß, S. C. Chapman, K. E. K. Coppin, et al., MNRAS 413, 2314 (2011), eprint 1012.0305.
- Torres (2004) D. F. Torres, Astrophys. J. 617, 966 (2004), eprint astro-ph/0407240.
- Ackermann et al. (2012) M. Ackermann, M. Ajello, A. Allafort, L. Baldini, J. Ballet, D. Bastieri, K. Bechtol, R. Bellazzini, B. Berenji, E. D. Bloom, et al., Astrophys. J. 755, 164 (2012), eprint 1206.1346.
- Lupton et al. (1999) R. H. Lupton, J. E. Gunn, and A. S. Szalay, Astron. J. 118, 1406 (1999), eprint astro-ph/9903081.
- Stoughton et al. (2002) C. Stoughton, R. H. Lupton, M. Bernardi, M. R. Blanton, S. Burles, F. J. Castander, A. J. Connolly, D. J. Eisenstein, J. A. Frieman, G. S. Hennessy, et al., Astron. J. 123, 485 (2002).

### v2

@@ 3 Results
Next, we applied the matching criterion defined in Equation 2 to our primary sample of 3,166 morphologically selected galaxy mergers identified from SDSS imaging. This resulted in identification of 38 spatial coincidences between Fermi 4FGL-DR4 sources and galaxy mergers from our sample. To assess the likelihood of true physical associations versus chance alignments, we computed Poisson based match probabilities for each pair using Eq. 2. Among these 38 matches, 21 exhibit p-values below 0.05, indicating a low probability of random association. We classify these as statistically secure matches. According to source classifications in the 4FGL-DR4 catalog, the 21 matches within the 4 ellipses include four Flat Spectrum Radio Quasars (FSRQs), six BL Lacertae objects (BLL), three Blazar Candidates of Uncertain type (BCUs), and three Radio Galaxies (RDGs). The remaining five are unassociated sources (UIDs) lacking confirmed multiwavelength counterparts. Notably, none of the secure matches is classified as a pulsar. The five UIDs span gamma-ray flux densities in the range , suggesting the possible existence of a previously unrecognized population of gamma-ray sources associated with galaxy mergers.
To investigate potential observational selection effects, we also compared the distributions of spectroscopic redshift () and -band magnitude for the full merger catalog and the subset of 21 statistically significant matched sources (). As shown in Figure 4, the redshift distribution of the matched mergers closely follows that of the entire catalog, with the majority of sources lying at . This indicates that the associations are not strongly biased toward unusually nearby systems. Next, we compared the -band magnitude distribution of the full merger sample consisting of 3,166 galaxies with that of the 21 matched mergers, as shown in Figure 5. The matched mergers are found to span magnitudes in the range –, corresponding to the brighter portion of the parent distribution. We also did additional tests to explore whether the quality of spatial associations depends on the optical brightness of the merger counterparts, we plot the angular separation, expressed in units of the Fermi source’s 95% positional uncertainty () against the SDSS -band magnitude of the matched mergers in Figure 6. We find no evidence for a correlation, indicating that the spatial consistency of the associations does not depend on the optical brightness of the mergers.
@@
Overall, our analysis reveals that approximately 0.67% (21 out of 3,166) of galaxy mergers exhibit statistically robust associations with gamma-ray sources. The predominance of blazar-type objects among these secure matches accounting for 62% is consistent with their known dominance in the extragalactic gamma-ray sky. However, the presence of radio galaxies and a significant fraction of unassociated sources (24%) suggests that merger-driven gamma-ray activity may originate from a broader range of physical mechanisms than previously recognized. To visualize these associations, Figure 7 presents a multi-panel view of Fermi sources with at least one galaxy merger candidate within the corresponding positional uncertainty ellipse. In each panel, the ellipse denotes the Fermi positional uncertainty region, the blue dot marks the Fermi source position, and the red cross indicates the galaxy merger counterpart with the lowest p-value. The plots are depicted in decreasing order of p-value. Moreover, in Table 1 we show all the secure matches with their coordinates, SDSS -mag, spectroscopic redshifts, flux densities, p-values, angular separation and classification. Uncertainties in -band magnitudes are from SDSS DR16 photometry; redshift uncertainties are from SIMBAD; flux errors are calculated from -band magnitudes. To explore marginal cases, we extended the analysis to include matches with , identifying 24 potential associations. Of these, 21 are already included in the secure sample (). The remaining three additional sources are 4FGL J1454.0+4927 (BCU), 4FGL J1523.2+2818 (BCU), and 4FGL J1223.3+1213 (BLL), with gamma-ray flux densities of , , and , and corresponding p-values of 0.069, 0.071, and 0.076, respectively. These sources are spatially associated with the SDSS galaxies SDSS J145506.73+493614.7, SDSS J152228.74+281402.9, and SDSS J122256.10+121739.2, which have spectroscopic redshifts of 0.0945, 0.0715, and 0.0905, respectively.
Furthermore, to assess the nature of the gamma-ray emission in the five unassociated Fermi sources in our sample, we examined their possible counterparts in the SDSS, the SIMBAD astronomical database (Wenger et al., 2000), the ROSAT All-Sky Survey (RASS) (Voges et al., 1999), and the Faint Images of the Radio Sky at Twenty Centimeters (FIRST) survey (Becker et al., 1994). Three sources appear to be associated with galaxies showing no evidence of AGN activity, while the remaining two show signs of contamination by background AGN or blazars (Massaro et al., 2009; D’Abrusco et al., 2019; Liu et al., 2011). 4FGL J0844.5+3035 is located 0.84 arcseconds from SDSS J084422.83+303556.3, a galaxy at redshift with no signs of AGN activity in its optical spectrum from SDSS DR16. The same galaxy is also listed as LEDA 1910153 in the Lyon-Meudon Extragalactic Database (LEDA), with a consistent redshift () (Zheng and Shen, 2020). It is undetected in both the RASS and FIRST surveys, disfavoring a background AGN or blazar scenario. Similarly, 4FGL J1112.0+2607 is offset by just 0.51 arcseconds from SDSS J111204.44+260240.8, a galaxy at redshift in SDSS DR16, with no optical signs and no detections in the RASS or FIRST catalogs (Duarte Puertas et al., 2017). Also, 4FGL J1600.4+0407 is just 0.14 arcseconds from SDSS J160049.74+041105.2, a galaxy at redshift with no AGN features in its SDSS spectrum and no detection in RASS or FIRST (Duarte Puertas et al., 2017). These cases represent plausible candidates for merger driven gamma-ray emission.
In contrast, the other two sources show evidence of AGN associations. 4FGL J1105.8+3944 lies only 0.15 arcseconds from SDSS J110553.81+394656.9, a galaxy at redshift with no AGN features in its SDSS DR16 optical spectrum (Toba et al., 2014; Kozieł-Wierzbowska et al., 2020). However, SIMBAD identifies a known BL Lac object, 87GB 110306.7+400253, at the same position (Massaro et al., 2009; D’Abrusco et al., 2019). 4FGL J1435.4+3338 is found within 1.0 arcsecond of 2MASS J14345466+3349199, which is classified in SIMBAD as an AGN with redshift (Liu et al., 2011). The nearest SDSS source, SDSS J143454.21+334934.5, is a Seyfert 2 galaxy at based on its optical spectrum (Toba et al., 2014; Gaia Collaboration, 2020). Future dedicated multiwavelength follow-up, particularly with high sensitivity instruments such as eROSITA (X-ray) or the Very Large Array (VLA) would be valuable to further confirm the absence of AGN signatures and establish the nature of the emission.
To evaluate the possibility of angular clustering among the 21 matched gamma-ray–merger pairs, we performed a two-point angular separation analysis. Our results indicate that the mean angular separation among all unique pairs is , with a median of . The mean nearest-neighbor distance is . Only 6 source pairs are within , and just one pair lies within . There are no pairs with separations less than . These results suggest mild angular clustering, but the effect is not strongly concentrated, and the majority of sources are widely dispersed across the sky. We conclude that the observed clustering is weak and unlikely to bias our match probability estimates, which are computed individually based on local error ellipse geometry. This analysis has been included in the manuscript for completeness.
To estimate the expected number of false associations under a null hypothesis, we constructed a control sample by spatially shifting the sky coordinates of all 3,166 merger candidates by random angular offsets within in both Right Ascension and Declination. This procedure preserves the sky distribution but breaks any physical associations between the mergers and Fermi sources. The shifted catalog was then cross-matched against the Fermi 4FGL-DR4 catalog using the same 4 error ellipse matching criterion employed for the real sample. In this shifted control sample, we found zero matches. In contrast, the original unshifted sample yielded 21 matches with -value . This result supports the interpretation that the observed matches are not random coincidences but instead indicate a statistically significant excess.
The predominance of blazar-type classifications (FSRQ, BLL) among our matched sources aligns with recent theoretical work suggesting that gamma-ray emission in merging SMBH systems arises from jet interactions and magnetohydrodynamic instabilities triggered by gravitational perturbations from secondary black holes 333Foord, A. (2024). A High-Energy Perspective on Merging SMBHs. 11th International Fermi Symposium. https://fermi.gsfc.nasa.gov/science/mtgs/symposia/eleventh/program/mon/Foord_AHighEnergyPerspectiveonMergingSMBHs.pdf.
4 Conclusions
@@
We have presented the first dedicated cross-matching analysis between galaxy merger catalogs and the Fermi-LAT 4FGL-DR4 gamma-ray source catalog (Ballet et al., 2023). From the first catalog, consisting of 70 galaxy pairs from the Canadian Network for Observational Cosmology Field Galaxy Redshift Survey (Patton et al., 2005), we found no statistically significant associations within the 4 positional uncertainty threshold (). The second catalog comprises a high-confidence sample of 3,166 morphologically selected galaxy mergers derived from the initial catalog of 328,151 candidate mergers from (Ackermann et al., 2018). Cross-matching with 4FGL-DR4 (Ballet et al., 2023) under a conservative 4 positional criterion yielded 38 spatial coincidences, of which 21 show p-values below 0.05. These associations are dominated by well-known gamma-ray source classes such as FSRQs, BL Lacs, and radio galaxies, lending support to the hypothesis that merger-driven processes can fuel AGN activity observable in gamma rays (Ellison et al., 2008, 2011; Scudder et al., 2012; Satyapal et al., 2014; Gabor et al., 2016; Kashiyama and Mészáros, 2014). Importantly, five secure associations correspond to unassociated Fermi sources, suggesting that merging systems may host previously unrecognized gamma-ray populations.
5 Discussion and Outlook
@@
While blazar-like AGN dominate the matched sample, the presence of radio galaxies, UIDs, and the broad range of observed gamma-ray flux densities point to a diversity of possible emission mechanisms. These may include merger-induced starbursts, large-scale shocks, or obscured AGN activity (Torres, 2004; Ackermann et al., 2012), underscoring the importance of deeper multi-wavelength follow-up campaigns. The recurrent appearance of unassociated sources within the secure sample highlights the potential of galaxy merger catalogs to serve as physically motivated priors for gamma-ray source identification.
As gamma-ray localization improves and forthcoming optical surveys such as the Vera Rubin LSST, Euclid, and DESI expand the census of galaxy mergers, the prospects for establishing robust associations should improve significantly. In addition, we plan to conduct a targeted search for specific merger candidates, such as UGC 12914/5, UGC 813/6, and VV 114, in a forthcoming work. Furthermore, future analyses could benefit from incorporating the future updated Fermi-LAT source catalogs as well as updated merger catalogs from ongoing and upcoming surveys, which will provide improved completeness and accuracy for cross-matching studies.
@@
| 4FGL J1606.5+2717 | 241.6173 | 27.1373 | 0.0480 | 3.70 | FSRQ | |||
| 4FGL J1245.8+0232 | 191.5785 | 2.6659 | 0.0344 | 3.31 | BLL | |||
| 4FGL J0815.6+3641 | 123.8059 | 36.7766 | 0.0252 | 3.77 | FSRQ | |||
| 4FGL J1518.6+0614 | 229.5286 | 6.3552 | 0.0274 | 2.68 | RDG | |||
| 4FGL J1112.0+2607 | 168.0185 | 26.0447 | 0.0182 | 2.64 | UID | |||
| 4FGL J1435.4+3338 | 218.7278 | 33.8222 | 0.0171 | 3.00 | UID | |||
| 4FGL J0813.4+2531 | 122.8163 | 25.1794 | 0.0113 | 3.99 | FSRQ | |||
| 4FGL J1008.0+0028 | 151.9620 | 0.5303 | 0.0111 | 3.13 | BLL | |||
| 4FGL J1145.50340 | 176.3724 | 3.6637 | 0.0122 | 0.83 | BLL | |||
| 4FGL J1600.4+0407 | 240.2073 | 4.1848 | 0.0137 | 3.42 | UID | |||
| 4FGL J1353.2+3740 | 208.3187 | 37.7076 | 0.0076 | 1.26 | BLL | |||
| 4FGL J0016.20016 | 3.9988 | 0.3049 | 0.0081 | 2.50 | FSRQ | |||
| 4FGL J1144.9+1937 | 176.2744 | 19.6365 | 0.0083 | 1.97 | RDG | |||
| 4FGL J1202.9+5141 | 180.7901 | 51.6782 | 0.0060 | 1.67 | BCU | |||
| 4FGL J1018.9+1043 | 154.6662 | 10.7793 | 0.0058 | 0.53 | BCU | |||
| 4FGL J0844.5+3035 | 131.0952 | 30.5990 | 0.0042 | 1.65 | UID | |||
| 4FGL J1441.7+1836 | 220.5586 | 18.6907 | 0.0035 | 3.91 | BCU | |||
| 4FGL J1411.8+5249 | 212.9420 | 52.7960 | 0.0028 | 2.64 | BLL | |||
| 4FGL J1105.8+3944 | 166.4743 | 39.7825 | 0.0026 | 2.26 | UID | |||
| 4FGL J1116.6+2915 | 169.1460 | 29.2497 | 0.0006 | 0.80 | RDG | |||
| 4FGL J1508.8+2708 | 227.1547 | 27.1589 | 0.0005 | 3.41 | BLL |
Acknowledgments
@@
SM thanks the Ministry of Education (MoE), Government of India, for their consistent support through the research fellowship, which has been instrumental in facilitating the successful completion of this work. The authors also thank Lizhong Zhang for useful correspondence. We also thank the anonymous referee for very useful and constructive feedback on our manuscript.